i appreciate all the help! thank you in advance!
(1 point) Archeologists have discovered a sample of ancient Egyptian skulls from 150 A.C. A sample of 5 Egyptian skull breadt 1

At least one of the answers above is NOT correct. (1 point) How much money do winners go home with from the television quiz s2

139.61 139.61 incorrect At least one of the answers above is NOT correct. (1 point) Archeologists have discovered a sample of3
Answer #1

8.1) let the date as, n = 5 132, 135, 137, 135, 135. Ex= 674 X = 134.8 8 {(xi-x² 3.2 12.8 4 S= 1.989 2=0.05 & df = n-1= 5+1=4c1= x I ta, df 95% c 1 for mean is. = 134.3 + 2.776 x 1,789 √5 = 13418 + 2.776 x 0.8 = 134.8 + 2.22 C 132.58, 137.02) 132.5894% CI for el is, x & te,df s 5551.93 31813, 8 I 2.046 × 15 31813.8 I 2.046 X 1433.5 31813.8 2933.19 ( 28880.61 34746.99) 7 299% CI for el is, X Į ta, da x S 126,33 1 4.03 X 8,07 126,33 € 4.03 X 3.29 126.33 + 13.26 ( 113.05 139.59 113.05 to 139.59